OpenStudy (mayankdevnani):

An echo repeats two syllables.If the velocity of sound is 330 m/s , then the distance of the reflecting surface is- a) 66.0 m b) 33.0 m c) 99.0 m d) 16.5 m

OpenStudy (mayankdevnani):

@amistre64 @TuringTest @hartnn @ganeshie8 @ghazi @theyatin

OpenStudy (ghazi):

i am not sure about this , but possibly its 66 m

OpenStudy (mayankdevnani):

yaa its right!!!

OpenStudy (mayankdevnani):

how do you do this? @ghazi

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (mayankdevnani):

no answer is wrong it gets 16.5m @theyatin

OpenStudy (mayankdevnani):

@ganeshie8 @geerky42 @Aperogalics

OpenStudy (anonymous):

oops then i must sleep. . .

OpenStudy (mayankdevnani):

oh!!!! koi baat nahi

OpenStudy (ghazi):

this will help you http://in.answers.yahoo.com/question/index?qid=20120713012737AAWgHSc

OpenStudy (anonymous):

can you explain the concept of syllable??

OpenStudy (mayankdevnani):

i don't know what is syllable??

OpenStudy (mayankdevnani):

@hartnn @ash2326 @jazy @agreene

OpenStudy (mayankdevnani):

@TuringTest

OpenStudy (anonymous):

well its very tough i see it as if minimum frequency gets us time t=0.05s minimum and as we have 2 syllables 0.05*2=0.1 ( as i have studied in link given by gazi) so the distance must be velocity* time=33meter but to strike wall and come back it becomes 66 as total. . . but here question is asking for the distance of surface so i am not quite sure. . .

OpenStudy (mayankdevnani):

@amistre64 @AccessDenied @ash2326 @dumbcow @hartnn @UnkleRhaukus @Callisto

OpenStudy (mayankdevnani):

@Shane_B @surdawi @AccessDenied @ajprincess @Yahoo!

OpenStudy (mayankdevnani):

@henpen

OpenStudy (amistre64):

im assuming the question is saying that you hear your echo 2 times

OpenStudy (anonymous):

You need the time between the 2 sounds

OpenStudy (amistre64):



OpenStudy (anonymous):

yey!!! that justifies my answer 16.5m which when multiplied by 4 gets an answer of 66m again explaining it. . . we can hear minimum 20hz that leads to T=0.05sec and thus the distance leads to velocity * time = 330 * 0.05=16.5 thus when multilied by 4 gets 66m. . .

OpenStudy (mayankdevnani):

why we multiplied it by 4 @theyatin @amistre64 @henpen

OpenStudy (amistre64):

the distance traveled between the "wall" and the source is quadrupled in order to hear the echo twice, as i pictured it up. assuming that im reading the question correctly to begin with

OpenStudy (mayankdevnani):

oh!!! i see

OpenStudy (amistre64):

imagine you and a friend throwing a ball back and forth; how far does the ball travel relative to your positions if you catch it 2 times and throw it 2 times?

OpenStudy (mayankdevnani):

i think 4 times. right

OpenStudy (amistre64):

correct

OpenStudy (mayankdevnani):

thnx... @amistre64

OpenStudy (amistre64):

hope its the right concept ;) good luck

OpenStudy (mayankdevnani):

what is syllables???

OpenStudy (amistre64):

im thinking that it pertains to the rebound that you hear. and not really referencing words perse

OpenStudy (mayankdevnani):

oh!!!

OpenStudy (amistre64):

if you stand in an echoy place and give a brisk woot!! you can hear how many times it echos back at you

OpenStudy (amistre64):

one of them old dead scientist guys used a long hallway that he knew the length of in order to estimate the speed of sound that way